OpenStudy (anonymous):

Solve the equation and check the solution: a/(a^2-36)+2/(a-6) = 1/(a+6) How should I start this?

random231 (random231):

okay : wat can u write a^2-36 as?

OpenStudy (anonymous):

it factors too (a-6) (a+6)

random231 (random231):

good now u can take (a-6) common!

random231 (random231):

are u getting it?

OpenStudy (anonymous):

\[\dfrac{a}{(a-6) (a+6)} + \dfrac{2}{(a-6)} = \dfrac{1}{(a+6)}\] What do I do from there?

random231 (random231):